# AI-Driven Mobility Management for High-Speed Railway Communications: Compressed Measurements and Proactive Handover

### v2

@@ I Introduction
High-speed railway (HSR) provides people with safe, fast, comfortable and economical modes of transportation, and has been widely developed around the world. Nevertheless, mobile communication systems are pivotal for ensuring rail safety, operations, maintenance, and delivering passenger information services. Emerging applications, such as railway multimedia dispatching and high-definition video monitoring on board and trackside, have posed stricter requirements on the reliability, data rate, and spectral efficiency of the next-generation, dedicated mobile communication systems for HSR. However, the high speed of trains entails formidable challenges due to the time-varying nature of wireless channels in HSR communications. These challenges include the increased signaling overhead and reduced system throughput.
Thanks to the abundant spectrum available at higher frequencies, communication systems using the millimeter wave (mmWave) technology avail of faster data transmission speeds and reduced latency compared to traditional low-frequency communications. For HSR communications, mmWave frequency bands can underpin increased data transmission rates that meet the demands of onboard applications and services, and provides low latency that is indispensable for real-time communication requirements. However, mmWave communication systems suffer from significant path loss, which restricts the effective range of transmission. To tackle this problem, beamforming with large-scale antenna arrays can be employed to generate finely directional beams, mitigating propagation loss and expanding signal coverage. However, the high mobility of trains entails pressing challenges in mobility management for HSR communications. While being connected, a train has to address two types of mobility, i.e., the beam-level and cell-level mobility as specified in the 3rd Generation Partnership Project (3GPP) Release 17 for fifth generation (5G) systems[1]. Beam-level mobility involves dynamically switching beams in which the train establishes a link path to a new beam within the same cell, ensuring communication with high performance. On the other hand, cell-level mobility focuses on cell handover between different cells. As trains move from one coverage zone to another, it is crucial to have a smooth cell handover to sustain continuous connectivity experience.
@@
Beam-level mobility management in 5G relies on periodic channel state information reference signals and synchronization signal blocks for beam sweeping, measurement, and reporting. However, frequent beam sweeping introduces significant latency and high overhead[2], making it impractical for HSR communications. Cell-level mobility management primarily relies on an event A3-based handover process, encompassing handover trigger, preparation, and execution. Before handover completion, the UE continues to communicate with the source base station (BS), often resulting in a sharp decline in the communication quality. Additionally, with a fixed handover period, the link quality between the high-speed terminal and source BS deteriorates faster, increasing the risk of handover failure. Hence, innovative handover mechanisms are urgently needed to address these limitations.
@@
In this article, we provide two AI-based beam-level and cell-level mobility management approaches for HSR communications. In Section II, we reveal the potential of AI-based spatial-temporal beam prediction for HSR communications. Particularly, by leveraging the sparsity and time-domain correlation of channels in HSR communications, we propose compressed spatial multi-beam measurements via compressive sensing (CS) to improve the spatial-temporal beam prediction accuracy in HSR communications. In Section III, we investigate AI-based cell-level mobility management for HSR scenarios and propose a novel AI-assisted cell handover mechanism, performing handover proactively before the signal quality deteriorates by AI-based handover event prediction. Section V introduces the challenges and opportunities encountered in AI-based mobility management for HSR communications, followed by our conclusions in Section VI. A summary of our overview of AI-based mobility enhancement schemes is given in Table I.
@@ II-A AI-Based Beam-Level Mobility Management in HSR Communications
When a user equipment (UE) performs a beam switch within the same cell, it initiates beam-level mobility. This process, which includes beam sweeping, beam measurement, beam reporting, and beam determination is managed within the cell, where the physical (PHY) and media access control (MAC) layers are sufficient to handle it, thereby avoiding the need for radio resource control (RRC) operations. The detailed beam-level mobility management process is shown in Fig. 2.
Beam-level mobility management enhancement involves three types of beam search methods: traditional methods including exhaustive search and hierarchical search, advanced numerical methods, and AI-based methods. The exhaustive beam search method examines all the candidate beams in the predefined codebook one by one to obtain beam quality information. The hierarchical beam search method provides a more efficient alternative, which first searches within a wide beam range and then further searches for a narrow beam within the angular range of the selected wide beam. However, traditional beam search algorithms suffer from a large measurement overhead, which is not suitable for HSR communications which require frequent beam switching. In addition, enhancement techniques, based on numerical methods have been proposed. More specifically, [4] exploited the tree search method for efficient beam training based on the hierarchical codebook, whilst [5] proposed an optimal beam training model based on Markov’s decision process. These numerical approaches mainly target general communication environments, making it challenging to tailor them for particular communication scenarios. In HSR communications, with some coarse-grained beam quality measurements of the serving cell, AI can be utilized to estimate the finely-grained optimal beam pair in the current instance or predict the optimal beam pair in future instances, which leads to reduced measurement overhead and enhanced throughput.
@@
Different tasks in beam-level mobility management require distinct inputs to the AI models in HSR communications. In order to instigate further discussion, there are two vital concepts for beam-level mobility management, namely Set A and Set B. [^2] https://www.3gpp.org/ftp/tsg_ran/WG1_RL1/TSGR1_109-e/Inbox/drafts/9.2.3.2/R1-220xxxx%20%5B109-e-R18-AIML-06%5D-Summary-1-v022-MediaTek_Intel.docx Set A comprises all narrow beams from the predefined codebook, while Set B encompasses the beams that have been measured and are utilized as inputs into the AI models. Various alternative model input methods exist with respect to the relationship between Set A and Set B. For example, the AI model could receive the input of RSRP measurements of partial narrow beams, making Set B a subset of Set A. This is the most commonly used method in HSR communications, which greatly reduces the measurement overhead. Another option is to utilize RSRP measurements of wide beams to estimate the quality of narrow beams, where Set B and Set A are distinct. In addition, whenever the RSRP measurements of all narrow beams are the inputs into the AI models, Set B can be the same as Set A. The first two types of inputs, where the inputs to the model are RSRP measurements of some narrow beams and all wide beams, are designed for temporal and spatial beam prediction models, aiming to reduce the measurement overhead by lowering the frequency of beam measurements, which makes them more suitable for latency-sensitive applications. In contrast, the last type of input relies on measuring the RSRP of all candidate beams to conduct temporal beam prediction, achieving the higher model prediction accuracy. This method is well-suited for applications that demand exceptionally high accuracy in prediction models and are less sensitive to latency. Instead of using RSRP, supplementary auxiliary information, like the position of the UE, its mobility vector, the signal-to-noise ratio (SNR), and the beam index can currently be utilized to provide broader spatial context to forecast the state of the link connection [13].
@@ II-B AI-Based Compressed Multi-Beam Measurement Scheme
In HSR communications, every train follows a predetermined path and performs the same routes periodically. Given sufficient training data, AI models are expected to perform exceptionally well in managing beam-level mobility owing to this intrinsic regularity. For instance, AI-based super-resolution techniques can be employed for spatial beam prediction to reduce the measurement overhead in HSR communications. Neural networks can receive inputs in the form of low-resolution beam measurements, consisting of partial narrow beams or wide beams, and output high-resolution spatial narrow beam predictions. The super-resolution-based spatial beam prediction can be implemented using a sub-pixel convolutional layer paired with an up-sampling layer [6], or a combination of CNN and FC [9].
@@
Here, we investigate DL-based spatial-temporal beam prediction using compressed multi-beam channel measurements. DL is employed to output the beam quality in the future slots using historical multi-beam channel measurements in the compressed domain, where the CS measurement matrix can be learned jointly. In the training phase, the complete beam channel information from historical instances serves as input into the model. The CS matrix serves as the first layer, namely the linear compression layer. The compressed channel information is fed into the remaining neural network to reconstruct the complete RSRP information of all candidate beams for a future instance. In the remaining module, a sub-pixel convolution layer and a 2D upsampling layer are applied before data is fed into the convolutional LSTM (Conv-LSTM) layers. These two layers learn to upsample low-resolution features from limited beam measurements, mapping them to high-resolution feature maps of the complete beam measurements. Then, the initial estimates of the complete beam measurement map are fed into the Conv-LSTM layers to calculate the temporal variations by leveraging historical data and local correlations. Finally, the network outputs a complete RSRP map for an upcoming time slot. As a learnable parameter, the CS measurement matrix is jointly optimized during the training of the model. The structure of the proposed AI model for beam-level mobility management, namely CSAI, is illustrated in Fig. 3(d).
In the inference phase, the trained measurement matrices are applied at the transceivers while the compressed channel measurements are fed into the trained neural network to conduct beam prediction. Meanwhile, the model can undergo periodic retraining or continuous updates through online learning. Detailed deployment information is shown in Fig. 2 Compared with the traditional sweeping method, the multi-beam compression offers to the beam prediction network richer information with the same number of beam measurements.
@@
Now, we compare the performance of AI-based spatial-temporal beam prediction methods, whose structures are shown in Fig. 3. The AI-based methods in Fig. 3(a) [8] and Fig. 3(b) [9] employ the traditional LSTM and CNN architecture, respectively, to predict the high-resolution beams. Compared with the CSAI in Fig. 3(d), the AI-based method in Fig. 3(c) [6] has the same network structure except for the beam measurement module. In the experiments, the channel used in our experiments is generated by QuaDRiGa, and we consider the UMa scenario for outdoor 30 GHz mmWave communications, while the UE follows a straight-line trajectory which aligns with the HSR scenario. Four distinct datasets, representing UE speeds of 60km/h, 120km/h, 350km/h and 500km/h, are utilized in the evaluation. These datasets comprise 40,000 training samples, 5,000 validation samples, and 5,000 test samples. The batch size is set to 256, and the training runs for a maximum of 800 epochs. The learning rate changes dynamically from $10^{-3}$ to $10^{-8}$, utilizing the adaptive moment estimation optimizer. We consider the prediction accuracy for the best beam with a compression ratio [^3] The compression ratio is defined as the ratio of the number of input beam measurements to the number of all beam measurements. of $\frac{1}{16}$. The simple non-AI approach straightforwardly uses the best one out of the partially measured present beams as the prediction. For three traditional AI-based beam prediction methods shown in Fig. 3(a) (b) (c), the model’s input uses equal interval sampling from predefined codebook to obtain beam measurement information with a compression ratio of $\frac{1}{16}$. As shown in Fig. 4, the AI-based spatial-temporal beam prediction methods outperform significantly the traditional non-AI method. As the speed of the movement increases, all methods generally show an anticipated degradation in the prediction accuracy. The CSAI demonstrates superior performance for all the cases. At the speed of 500km/h, the CSAI is still able to yield a prediction accuracy of 83.76%. The Conv-LSTM network shown in Fig. 3(c) requires 3.84M flops and 0.53M parameters, whereas the proposed CSAI model utilizes 3.88M flops and 0.136M parameters. This demonstrates that the proposed model improves prediction performance with a negligible increase in computational cost compared to the fixed down-sampling beam pattern approach.
@@ III-A AI-Based Cell-Level Mobility Management in HSR Communications
Cell-level mobility management deals primarily with cell handovers, initiating the transition of a user’s session to a new cell when the signal strength from the serving cell falls below a predefined threshold. Given the rapid movement in HSR communications, cell handovers occur frequently, and failures in these handovers would unavoidably compromise the user experience. Handover decision is performed by the source cell based on UE measurement. UE measurement should be met by certain criteria, for example, event A3 which will be triggered when the difference between the layer 3 RSRP (L3-RSRP) [^4] L3-RSRP represents the cell-level RSRP, obtained by averaging or maximizing the RSRP of all beam pairs in each cell. measurement of the target cell and the L3-RSRP measurement of the source cell exceeds a predefined offset. If the above criteria are met during the time to trigger (TTT) duration, the UE reports the measurements and the source cell initiates the handover preparation phase. Afterwards, the cell handover procedure initiates the execution after the UE receives the handover command message, whose task is to provide RRC reconfiguration message to the target cell. The specific cell handover process is shown in Fig. 5.
In state 1, the UE measures the beam quality and sends the measurement report to the source BS. In this state, the UE remains connected to the source cell. When the mobility of the UE is low or the UE is moving within a macro cell, the traditional handover mechanism can sustain good performance. However, when the mobility of UE is high or the UE is moving across micro cells of high density, the traditional handover mechanism would result in unintended events, e.g., handover failure, radio link failure (RLF), ping-pong phenomenon, throughput loss. In HSR communications, the trains move at high speed, resulting in a sharp decline of the communication quality of the link between the UE and the source BS during state 1. Leveraging its robust environmental awareness and learning abilities, AI/ML technology can address the aforementioned issues.
@@
For the HSR scenario, we also explore various model inputs and outputs tailored to its specific requirements. One input option for AI models is the L3-RSRP of each cell, with the output being either the L3-RSRP of each cell or the L3-RSRP of the top $K$ cells. The dimension of L3-RSRP information is lower than that of layer 1 RSRP (L1-RSRP) [^5] L1-RSRP represents the beam-level RSRP, which is the RSRP of each beam pair of each cell. However, acquiring L3-RSRP, which requires averaging data over multiple time slots, introduces extra latency which is not desired in HSR communications. Another viable approach involves using L1-RSRP as both the input to and the output of the AI models. The AI model output is then filtered to derive the L3-RSRP, which is used to execute cell handovers. For this option, the output L1-RSRP of the AI model can be applied to both beam-level and cell-level mobility management. Nevertheless, L1-RSRP encompasses extensive information that exceeds the requirements for handover, while the correlation between the input and output L1-RSRP is intricate, thereby amplifying the learning challenge for the AI models. Moreover, leveraging L1-RSRP to forecast L3-RSRP via AI can be an advantageous strategy. This method reaps the benefits of the above two approaches, making use of the abundant data in RSRP to enhance the AI model’s prediction accuracy, while also streamlining the mapping between the model’s input and output.
@@
Our CNN/LSTM Compressed RSRP images RSRP of all beams of all cells Using two AI models to optimize the shortcomings of the handover mechanism. RLF rates

### v3

@@ I Introduction
High-speed railway (HSR) provides people with safe, fast, comfortable and economical modes of transportation, and has been widely developed around the world. Nevertheless, mobile communication systems are pivotal for ensuring rail safety, operations, maintenance, and delivering passenger information services. However, the high speed of trains entails formidable challenges due to the time-varying nature of wireless channels in HSR communications. These challenges include the increased signaling overhead and reduced system throughput.
Thanks to the abundant spectrum available at higher frequencies, communication systems using the millimeter wave (mmWave) technology avail of faster data transmission speeds and reduced latency compared to traditional low-frequency communications. For HSR communications, mmWave frequency bands can underpin increased data transmission rates that meet the demands of onboard applications and services, and provides low latency that is indispensable for real-time communication requirements. However, mmWave communication systems suffer from significant path loss, which restricts the effective range of transmission. To tackle this problem, beamforming with large-scale antenna arrays can be employed to generate finely directional beams, mitigating propagation loss and expanding signal coverage. However, the high mobility of trains entails pressing challenges in mobility management for HSR communications. While being connected, a train has to address two types of mobility, i.e., the beam-level and cell-level mobility as specified in the 3rd Generation Partnership Project (3GPP) Release 17 for fifth generation (5G) systems[1]. Beam-level mobility involves dynamically switching beams during which the train establishes a link path to a new beam within the same cell, ensuring communication with high performance. On the other hand, cell-level mobility focuses on cell handover between different cells. As trains move from one coverage zone to another, it is crucial to have a smooth cell handover to sustain a continuous connectivity experience (Figure 1).
@@
Beam-level mobility management in 5G relies on periodic channel state information reference signals and synchronization signal blocks for beam sweeping, measurement, and reporting. However, frequent beam sweeping introduces significant latency and high overhead[2], making it impractical for HSR communications. Cell-level mobility management primarily relies on an event A3-based handover process, encompassing handover trigger, preparation, and execution. Before handover completion, the UE continues to communicate with the source base station (BS), often resulting in a sharp deterioration in the communication quality. Additionally, with a fixed handover period, the link quality between the high-speed terminal and source BS deteriorates faster, increasing the risk of handover failure. Hence, innovative handover mechanisms are urgently needed to address these limitations.
@@
In this article, we provide two AI-based beam-level and cell-level mobility management approaches for HSR communications. Particularly, by leveraging the sparsity and time-domain correlation of channels in HSR communications, we propose compressed spatial multi-beam measurements via compressive sensing (CS) to improve the spatial-temporal beam prediction accuracy in HSR communications. We also investigate AI-based cell-level mobility management for HSR scenarios and propose a novel AI-assisted cell handover mechanism, performing handover proactively before the signal quality deteriorates by AI-based handover event prediction. A summary of our overview of AI-based mobility enhancement schemes is given in Table I.
@@ II-A AI-Based Beam-Level Mobility Management in HSR Communications
When a user equipment (UE) performs a beam switch within the same cell, it initiates beam-level mobility. This process, which includes beam sweeping, beam measurement, beam reporting, and beam determination is managed within the cell, where the physical and media access control layers are sufficient to handle it, thereby avoiding the need for radio resource control (RRC) operations. The detailed beam-level mobility management process is shown in Fig. 2. The receiver sequentially scans all predefined beam pairs and evaluates their quality. The beam pair with the highest RSRP is then chosen for beam switching.
Beam-level mobility management enhancement involves three types of beam search methods: traditional methods including exhaustive search and hierarchical search, advanced numerical methods, and AI-based methods. The exhaustive beam search method examines all the candidate beams in the predefined codebook one by one to obtain the beam quality information. The hierarchical beam search method provides a more efficient alternative, which first searches within a wide beam range and then further searches for a narrow beam within the angular range of the selected wide beam. However, traditional methods suffer from high measurement overhead, making them unsuitable for HSR communications, which require frequent beam switching. Numerical methods have also been proposed to address this challenge. For example, [4] employed tree search for efficient beam training based on a hierarchical codebook, while [5] proposed an optimal beam training model using Markov decision processes. These numerical approaches mainly target general communication environments, making it challenging to tailor them for particular communication scenarios. In HSR communications, with some coarse-grained beam quality measurements of the serving cell, AI can be utilized to estimate the finely-grained optimal beam pair in the current instance or predict the optimal beam pair in future instances, which leads to reduced measurement overhead and enhanced throughput.
@@
[…]
@@ II-B AI-Based Compressed Multi-Beam Measurement Scheme
In HSR communications, every train follows a predetermined path and performs the same routes periodically. Given sufficient training data, AI models are expected to perform exceptionally well in managing beam-level mobility owing to this intrinsic regularity. For instance, the super-resolution-based spatial beam prediction can be implemented using a sub-pixel convolutional layer paired with an up-sampling layer [6], or a combination of CNN and FC [9].
@@
Here, we investigate DL-based spatial-temporal beam prediction using compressed multi-beam channel measurements. DL is employed to output the beam quality in the future slots using historical multi-beam channel measurements in the compressed domain, where the CS measurement matrix can be learned jointly. A general model is initially trained and deployed at each cell, where it can undergo periodic retraining or continuous updates using locally collected data. In the training phase, the complete beam channel information from historical instances serves as input into the model. The CS matrix serves as the first layer, namely the linear compression layer. The compressed channel information is fed into the remaining neural network to reconstruct the complete RSRP information of all candidate beams for a future instance. In the remaining module, the compressed features are first processed by a 2D convolution (Conv-2D) layer. This layer extracts spatial features from the compressed representation, ensuring that essential beam information is preserved before further processing. Subsequently, a sub-pixel convolution layer and a 2D upsampling layer are applied before data is fed into the convolutional LSTM (Conv-LSTM) layers. These two layers learn to upsample low-resolution features from limited beam measurements, mapping them to high-resolution feature maps of the complete beam measurements. Then, the initial estimates of the complete beam measurement map are fed into the Conv-LSTM layers to calculate the temporal variations by leveraging historical data and local correlations. Finally, the network outputs a complete RSRP map for an upcoming time slot. As a learnable parameter, the CS measurement matrix is jointly optimized during the training of the model. The structure of the proposed AI model for beam-level mobility management, namely CSAI, is illustrated in Fig. 3(d). In the inference phase, the trained measurement matrices are applied at the transceivers while the compressed channel measurements are fed into the trained neural network to conduct beam prediction.
@@
Now, we compare the performance of AI-based spatial-temporal beam prediction methods, whose structures are shown in Fig. 3. Figure 3(a) [8] and Fig. 3(b) [9] employ the traditional LSTM for temporal modeling and CNN for spatial feature extraction, respectively, to predict the high-resolution beams. The CSAI model in Fig. 3(d) and the AI-based method in Fig. 3(c) [6] share the same Conv-LSTM architecture which integrates convolutional operations into the LSTM framework to jointly capture the spatial and temporal features. The only distinction between the two lies in the design of the beam measurement module. In the experiments, the channel is generated using a system-level simulation platform which we build based on QuaDRiGa. We deploy 7 BSs, where each BS is equipped with three half-wavelength spaced uniform plane arrays (UPAs) with a sector orientation of 30, 150 and -90 degrees. We further consider the urban macro scenario for outdoor 30 GHz mmWave communications, while the UE follows a straight-line trajectory which aligns with the HSR scenario. Four distinct datasets, representing UE speeds of 60km/h, 120km/h, 350km/h and 500km/h, are utilized in the evaluation. These datasets comprise 40,000 training samples, 5,000 validation samples, and 5,000 test samples. The batch size is set to 256, and the training runs for a maximum of 800 epochs. The learning rate changes dynamically from $10^{-3}$ to $10^{-8}$, utilizing the adaptive moment estimation optimizer. We consider the prediction accuracy for the best beam with a compression ratio [^3] The compression ratio is defined as the ratio of the number of input beam measurements to the number of all beam measurements. of $\frac{1}{16}$. The simple non-AI approach straightforwardly uses the best one out of the partially measured present beams as the prediction. For the three traditional AI-based beam prediction methods shown in Fig. 3(a) (b) (c), the model’s input uses equal interval sampling from predefined codebook to obtain beam measurement information with a compression ratio of $\frac{1}{16}$. As shown in Fig. 4, the AI-based spatial-temporal beam prediction methods outperform significantly the traditional non-AI method. As the speed of the movement increases, all methods generally show an anticipated degradation in the prediction accuracy. The CSAI demonstrates superior performance for all the cases. At the speed of 500km/h, the CSAI is still able to yield a prediction accuracy of 83.76%. The network in Fig. 3(a) requires 0.145M flops and 0.048M parameters, while the network in Fig. 3(b) has 0.146M flops and 0.104M parameters. The Conv-LSTM network in Fig. 3(c) demands 3.84M flops and 0.53M parameters. In comparison, the proposed CSAI model uses 3.88M flops and 0.136M parameters, showing that it achieves superior prediction performance with only a minimal increase in computational cost compared to the Conv-LSTM network, which uses a fixed down-sampling beam pattern approach.
@@ III-A AI-Based Cell-Level Mobility Management in HSR Communications
Cell-level mobility management deals primarily with cell handovers, initiating the transition of a user’s session to a new cell when the signal strength from the serving cell falls below a predefined threshold. Given the rapid movement in HSR communications, cell handovers occur frequently, and failures in these handovers would unavoidably compromise the user experience. Handover decision is performed by the source cell based on UE measurements. UE measurements should meet by certain criteria, for example, event A3 which will be triggered when the difference between the layer 3 RSRP (L3-RSRP) [^4] L3-RSRP represents the cell-level RSRP, obtained by averaging or maximizing the RSRP of all beam pairs in each cell. measurement of the target cell and the L3-RSRP measurement of the source cell exceeds a predefined offset. If the above criteria are met during the time to trigger (TTT) duration, the UE reports the measurements and the source cell initiates the handover preparation phase. Afterwards, the cell handover procedure initiates the execution after the UE receives the handover command message, whose task is to provide RRC reconfiguration message to the target cell. The specific cell handover process is shown in Fig. 5.
During state 1, the UE measures beam quality and reports it to the source BS while remaining connected to the source cell. In HSR communications, high-speed movement leads to a rapid decline in link quality during this stage, often resulting in issues such as handover failure, radio link failure (RLF), ping-pong effects, and throughput degradation. Leveraging its strong environmental awareness and learning capabilities, AI/ML presents a promising solution to mitigate these challenges by predicting the RSRP during this state.
@@
For the HSR scenario, we explore various model inputs and outputs tailored to its specific requirements. One option is using the L3-RSRP of each cell as input, with the output being either the L3-RSRP of each cell or the top $K$ cells. L3-RSRP has lower dimensionality than L1-RSRP. [^5] L1-RSRP represents the beam-level RSRP, which is the RSRP of each beam pair of each cell. However, calculating L3-RSRP, which involves averaging over multiple time slots, introduces additional latency, undesirable in HSR communications. Another viable approach is to use L1-RSRP as both the input and output of the AI model. The output L1-RSRP can then be processed to derive the L3-RSRP for cell handovers. This method allows L1-RSRP to be used for both beam-level and cell-level mobility management. However, L1-RSRP contains more information than necessary for handovers, and its complex correlation between input and output increases the learning challenge. Moreover, leveraging L1-RSRP to forecast L3-RSRP via AI can be an advantageous strategy. This method reaps the benefits of the above two approaches, making use of the abundant data in RSRP to enhance the AI model’s prediction accuracy, while also streamlining the mapping between the model’s input and output.
@@
Our CNN/LSTM Compressed RSRP images RSRP of all cells Using two AI models to optimize the shortcomings of the handover mechanism. RLF rates